OpenStudy (anonymous):

If (x + yi) + (2(x + yi) + 3i) = 9, what is x + yi?

OpenStudy (anonymous):

open the bracket, what do you have?

OpenStudy (anonymous):

IM LOST

OpenStudy (anonymous):

(x+yi) + 2(x+yi) +3i =9, right?

OpenStudy (anonymous):

distribute the 2 in x + yi

OpenStudy (anonymous):

now, if 1(cstitt) + 2 (cstitt) =? (cstitt)?? or 1cookie + 2 cookies = ? cookies?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

SO ANSWER WILL BE 3+i?

OpenStudy (anonymous):

yes!! so the first 2 terms is 3(x+yi) , the last term from the left hand side is till 3i total, you have 3(x+yi) +3i =9 got me so far?

OpenStudy (anonymous):

yep thaank youu

OpenStudy (anonymous):

now, divide both sides by3,you get (x+yi) +i =3 ok?

OpenStudy (anonymous):

minus i (-i) both sides, you have the answer for (x+yi) = 3-i